Status Report

NOAA SATOPS Morning Report: Tuesday, June 26, 2012

By SpaceRef Editor
June 26, 2012
Filed under ,

SATOPS Morning Report: Tuesday, June 26, 2012

Customers or users who want to call to ask about problems or report outages can use the following contacts:
– For questions about data quality (noisy data, dropouts, schedules), GOES or POES products, or status of spacecraft can call the SAB Shift Supervisor at 301-763-8444
– To report an outage or document a problem, call the ESPC Operations Crew Lead at 301-457-5218. All LRIT data users are encouraged to report problems to: [email protected]

DOY 178

Critical Weather Day declared June 23, 2100Z; expected end June 30, 0000Z for Tropical Storm Debby..

GOES
All GOES operations were nominal over the past 24 hours.
No anomalies affecting product delivery via GVAR or s/c health & safety.
For GOES status and other information please see:
http://www.oso.noaa.gov/goes

POES
All POES operations were nominal over the past 24 hours.
No anomalies affecting product delivery or s/c health & safety.
For a detailed status and description of POES see:
http://www.oso.noaa.gov/poes

DMSP
DOY 178 26 June 2012 00:28Z: Ops-53 CAT 2 data loss declared at McMurdo; Spacecraft Seconds—604, Playback Seconds—151. Raytheon software fix to be deployed today to fix underlying cause.

All other DMSP operations were nominal over the past 24 hours.
No anomalies affecting product delivery or s/c health & safety.
For more DMSP Information please see:
http://dmsp.ngdc.noaa.gov/dmsp.html

JASON-2
All JASON-2 operations were nominal over the past 24 hours.
No anomalies affecting product delivery or s/c health & safety.

ESPC Reported data outage and delays
DOY 175 23 June 2012 0342Z – DOY 178 26 June 2012 0723Z: ESPC did not receive Cape Ferguson HRPT data. Data flow has been restored.

Questions or comments to [email protected]

SpaceRef staff editor.